is this statement true? explain.


distance travelled by a vertically projected body in the last second of its upward journey will be same as the distance travelled by a freely falling body in the first second.Is this independent of initial velocity of projection?

Grade:

2 Answers

RAJORSHI PAUL
12 years ago

Tine taken to reach the topmost point=u/g

So distance covered in (u/g)th sec=u-1/2g(2u/g-1)=1/2g which is indept of u

Now distance covered by a freely falling body in 1st sec=1/2gx1=1/2g

so the distances are same

basit ali
12 years ago

this statement is true. yes this is independent of initial velocity of projection. because when body goes to upward then at max.height its velocity becomes zero. and only the 'g' acts downward.

 if u like this plz approve it!!!Smile